OpenStudy (empty):

Can a rational number (not including 0, 1, or -1) be raised to an irrational power be rational?

ganeshie8 (ganeshie8):

I think I have asked this question sometime back \[(a/b)^x = p/q \iff x = \log_{a/b} (p/q) \]

ganeshie8 (ganeshie8):

There exist irrational numbers that are logarithm of a rational number relative to a rational number base. So we're done.

OpenStudy (empty):

Awesome thanks that makes a lot of sense!